OpenStudy (anonymous):

What is the equation of the line that passes through (0, 3) and (7, 0)? A. y=3/7x-3 B. y=3/7x+3 C. y=-3/7x+3 D. y=-3/7x-3

OpenStudy (abdullah1995):

Hi, what have you done so far?

OpenStudy (anonymous):

nothing

OpenStudy (anonymous):

ok so i (0,3) whick number is your x??

OpenStudy (anonymous):

0

OpenStudy (anonymous):

ok so you y is what?

pooja195 (pooja195):

\[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

3

pooja195 (pooja195):

Lets find the slope first

pooja195 (pooja195):

plug in the numbers

OpenStudy (anonymous):

ok good so we know its not c or d becaus y = 3

OpenStudy (anonymous):

3-0=3 0-7=-7

OpenStudy (anonymous):

ok so out of a and b do u have an idea of which one youthink it is?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

yea thats what i think :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome :)